OpenStudy (anonymous):

What are the classifications of each system? Column A Column B 1. 2. 3. 4. A. consistent independent B. coincident C. inconsistent

OpenStudy (anonymous):

so # 1: 4X-2Y= 10 2X-Y=5 # 2: Y=2X-3 -2X+Y= -5 #3: 2X-5Y= 14 3X+4Y= 10 #4: Y=4X-3 12X-3Y= 9

OpenStudy (anonymous):

that is column A ^

OpenStudy (anonymous):

A. consistent independent B. coincident C. inconsistent

OpenStudy (anonymous):

that is column B ^

OpenStudy (airjordan200):

yep B

OpenStudy (anonymous):

all of them r not b though i have to see what each one is they r all different

OpenStudy (airjordan200):

ik i mean number 2 is b

OpenStudy (airjordan200):

#1 A

OpenStudy (anonymous):

HURRY UP !!!!!!!

OpenStudy (airjordan200):

there is only 2 i see

OpenStudy (anonymous):

4 don't u see 4 problems

OpenStudy (airjordan200):

nnooooo..............

OpenStudy (anonymous):

yes dude i am looking at them right now

OpenStudy (airjordan200):

not me my computr might be glitching

OpenStudy (anonymous):

# 3: 2X-5Y=14 3X+4Y=10

OpenStudy (anonymous):

that is # 3

OpenStudy (anonymous):

# 3: Y=4X-3 12X-3Y=9

OpenStudy (anonymous):

that is # 4

OpenStudy (airjordan200):

number 4 is C

OpenStudy (airjordan200):

which leaves 3 with A